OpenStudy (anonymous):

if dy/dx=4y and y=4 when x=0, then y= a)4e^4x b)e^4x c)3+e^4x d)4+e^4x e)2x^2+4

OpenStudy (ash2326):

\[\frac{dy}{dx}=4y\] It can be written as \[\frac{dy}{y}=4dx\] Now integrate both sides \[\int \frac{dy}{y}=\int 4 dx\] Can you do this ?

OpenStudy (anonymous):

d.

OpenStudy (ash2326):

How did you get this?

OpenStudy (anonymous):

i got \[\int\limits_{?}^{?}1/y* dy= 4x+c\]

OpenStudy (anonymous):

then ln(y)=4x+c ln(4)=4(0)+c ln4=c

OpenStudy (ash2326):

Doing great:)

OpenStudy (anonymous):

then e^ln(y)=e^(4x+ln4)

OpenStudy (anonymous):

y=e^4x+e^lne =e^4x+4

OpenStudy (anonymous):

opps, y=e^4x+e^ln4

OpenStudy (ash2326):

Just one mistake \[\large e^{\ln y}=e^{4x+\ln 4}\] When two same number with different powers are multiplied, powers add so \[\large y= e^{4x}\times e^{\ln 4}\] \[\large y=4e^4x\]

OpenStudy (anonymous):

but the e and ln cancel, dont they?

OpenStudy (ash2326):

Yeah they cancel

OpenStudy (ash2326):

\[\huge e^{\ln 4}=4\]

OpenStudy (ash2326):

Do you understand? this is also like that \[2^{x+y}=2^x \times 2^{y}\]

OpenStudy (anonymous):

not really...

OpenStudy (ash2326):

Where you have doubt? \[\large e^{4x+\ln 4}= e^{4x} \times e^{\ln 4}\] Do you get this?

OpenStudy (anonymous):

my teacher taught me to break it up, distribute the e through

OpenStudy (anonymous):

ohh and when i do that change the sign to multiplication

OpenStudy (ash2326):

Yeah , it changes to multiplication just like \[2^4=2^{2+2}=2^2 \times 2^2\]

OpenStudy (anonymous):

ohh okay. got it. thanks for your patience

OpenStudy (ash2326):

No worries:D Glad to help:P